A crystal has a=2,b=3 and c=3oA and α=β=γ=90o. What is the length of the longest diagonal?
1
Expert's answer
2020-03-24T14:01:22-0400

Unit cell is cuboid. The longest diagonal is the volume diagonal, its length is:

L = sqrt(a2+b2+c2) = sqrt(4+9+9) = sqrt(22) = 4.69 A